OpenStudy (robot):

A snowmobile travels 32 miles in 213 hours. To the nearest tenth, what was the average speed of the snowmobile?

OpenStudy (google):

wow i could walk further in that amount of time

OpenStudy (robot):

2 1/3

OpenStudy (google):

its 2 1/3? or 213

OpenStudy (robot):

2 1/3

OpenStudy (mathstudent55):

The average speed is the unit rate. You need miles per hour, which means the number of miles in 1 hour. \(\dfrac{32~miles}{213~hours}\) Since you want the denominator to be 1 hour, and it is 213 hours, divide the numerator and denominator by 213. \(\dfrac{32\div 213~miles}{213\div 213~hours}\) This will give you 1 hour in the denominator which is what miles per hour means.

OpenStudy (google):

its 2 1/3 nots 213

OpenStudy (google):

ill help you

OpenStudy (robot):

ok thx

OpenStudy (google):

Hold on 1 second

OpenStudy (robot):

ok

OpenStudy (mathstudent55):

The idea is still the same. Divide the distance by the time. \(\dfrac{32~miles}{2 \frac{1}{3} ~hours}\)

OpenStudy (google):

1/3= 0.33 2 1/3 = 2.3333 32/2.33 = 13.7339056 Nearest tenth = 13.73

OpenStudy (google):

i think, idk for sure

OpenStudy (google):

or 13.7

OpenStudy (mathstudent55):

\(=\dfrac{32~miles}{1} \div 2 \dfrac{1}{3} ~hours \) \(=\dfrac{32~miles}{1} \div \dfrac{7~hours}{3} \) \(=\dfrac{32~miles}{1} \times \dfrac{3}{7 ~hours} \) \(=\dfrac{32 \times 3~miles}{1 \times 7~hours} \) \(= \dfrac{96~miles}{7~hours} \) \(= 13.714 ~mph\) Now round off to the nearest tenth.

OpenStudy (google):

i think its 13.7

OpenStudy (robot):

thx

OpenStudy (google):

np

OpenStudy (google):

fan me if you wanna know when im online and if you need more help